OpenStudy (anonymous):

Which statement is true? In any right triangle, the sine of one acute angle is equal to the cosine of the other acute angle. In any right triangle, the sine of one acute angle is equal to the sine of its complementary angle. In any right triangle, the cosine of one acute angle is equal to the cosine of its complementary angle. In any right triangle, the sum of the sine of one acute angle and the cosine of the other acute angle is 1.

OpenStudy (michele_laino):

I think that only the first statement is true

OpenStudy (anonymous):

Yeah that's right :) Thank you!

OpenStudy (michele_laino):

in order to vrify tha, we can wrie: \[\cos(90-\alpha)=\cos 90 \cos \alpha+\sin 90 \sin \alpha=\sin \alpha\]

OpenStudy (michele_laino):

thank you!